faced with rising fax costs a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. the firm examined 35 randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages. (a) at the .01 level of significance, is the true mean greater than 10? (b) use excel to find the right-tail p-value.
Answer by stanbon(57984)

Faced with rising fax costs, a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. The firm examined 35 randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages.
(a) At the .01 level of significance, is the true mean greater than 10?
Ho: mu = 10
Ha: mu > 10
---------------
Critical value for one-tail test with alpha=1% and df=34 is 2.3263
Test statistic:
t(14.44) = (14.44-10)/[4.45/(sqrt35)] = 5.9028
Conclusion: Since test stat is greater than critical value, reject Ho.
The test provides statistical evidence that mu > 10
--------------------------------------------------
(b) Use Excel to find the right-tail p-value.
p-value = P(5.9028 < t < 10) = 0.00000057579...
Since the p-value is so much smaller than 1% this is very stong evidence for rejecting Ho and supporting Ha.
=========================
Cheers,
Stan H.